OpenStudy (anonymous):

Specify the domain of the function. y=1/2^x-1

OpenStudy (anonymous):

Okay multiply by 2, what do you get?

OpenStudy (anonymous):

Should be 2y = 2/2^x-1

OpenStudy (anonymous):

2 is the same as 2^1 right?

OpenStudy (anonymous):

Now you have the same base for your numerator and denominator

OpenStudy (anonymous):

So, business as usual 2^1/2^x-1 = 2^1-(x-1). Remember to distribute the negative.

OpenStudy (anonymous):

why would i multiply by 2 ?

OpenStudy (anonymous):

sorry the -1 is not with the x

OpenStudy (anonymous):

it is 2^x -1

OpenStudy (anonymous):

so its y = (.5^x) - 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To give your numerator the same base as your denominator - so that you can use your subtraction rule.

OpenStudy (anonymous):

You use that subtraction rule to get it out of the denominator, so that you no longer have a fraction.

OpenStudy (anonymous):

Okay yea that changed it lol, one sec

OpenStudy (anonymous):

i know that it is not all real numbers

OpenStudy (anonymous):

is it negative infinite to 0 then 0 to infinite

OpenStudy (anonymous):

not including zero

OpenStudy (anonymous):

Yes, otherwise whatever the base is would be 1^10 instead of .5^x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Okay got it.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y= (1/2)^x -1 y-1 = (1/2)^x y-1 = 1^x/2^x (1 to any power is still 1 - 1x1x1=1) y-1 = 1/2^x y-1 = 2^-x Now we take the log (base of the exponent) of the other side of the equation. This takes the x out of the exponent. -x = log<base2>(y-1) x = -log<base2>(y-1)

OpenStudy (anonymous):

Magic!

OpenStudy (anonymous):

thanks i was just looking for the interval notation